A spacecraft passes you at 0.29 the speed of light. By what factor would its relativistic energy increase if its speed doubled?
Answer #1

using the equation E = m / ( 1 - V2 / C2 )1/2

E2 / E1 = ( 1 - V12 / C2 )1/2 / ( 1 - V22 / C2 )1/2  

E2 / E1 = ( 1 - 0.292 X C2 / C2 )1/2 / ( 1 - 2 X 0.292 C2 / C2 )1/2  

E2 / E1 = ( 1 - 0.292 )1/2 / ( 1 - 2 X 0.292 )1/2

E2 / E1 = 1.1748